[figure: 9.7 cm]

DETERMINATION OF THE INTERFACIAL

MAGNETO-OPTICAL PROPERTIES IN SPUTTERED Fe/Au

MULTILAYER STRUCTURES

L. Uba^{a}, A. Polewko^{a}, S. Uba^{a}, R. Gontarz^{b}, A.N. Yaresko^{c}, V.N. Antonov^{d} aInstitute of Physics, University of Białystok, Lipowa 41, PL-15-424 Białystok, Poland b Institute of Molecular Physics, PAS, Smoluchowskiego 17, PL-60-179 Poznań, Poland cMax-Planck Institute for the Chemical Physics of Solids, D-01187 Dresden, Germany

dInstitute of Metal Physics, 36 Vernadskii Street, 252142 Kiev, Ukraine

The Fe/Au multilayers (MLS) prepared by dc- sputtering on GaAs(001) and Si(100) substrates with 60 nm Au buffer layer have been studied by spectroscopic ellipsometry and Kerr effects in polar and longitudinal geometry at the photon energies ranging from 0.8 to 5.8 eV. From the x-ray diffraction analysis it follows that the Fe/Au films studied exhibit well defined layered structure with the fcc crystal structure and pronounced (111) texture. The effective dielectric tensor components of the fcc Fe/Au MLS structures were derived from the experimental ellipsometry and magneto-optical (MO) spectra. In order to obtain a quantitative understanding of the MO response of the Fe/Au MLS we have performed extended multiple reflection calculations as well as calculations based on the band structure LMTO calculations for various models of the MLS structure. The results obtained show a crucial role of the interface volume and its structure in the formation of the MO response of the Fe/Au MLS. The differences between the MO properties measured for in-plane and out-of plane sample magnetization direction, firstly observed in (001) Fe/Au superlattices, are established for sputter deposited (111) Fe/Au studied and its relation with interface structure has been examined.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : L. Uba

Address for correspondence : Institute of Experimental Physics University of Białystok Lipowa 41 PL-15-424 Białystok Poland

Email address : ubaluba@venus.uwb.edu.pl; ubaluba@alpha.uwb.edu.pl